OpenStudy (anonymous):

Triangle STU is similar to triangle VWX. Which statement is true about the two triangles? Angles S and V are congruent. Angles T and U are congruent. Angles T and W are proportional. Angles V and X are proportional.

OpenStudy (anonymous):

@Anaise

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@happy_to_help

OpenStudy (anonymous):

plz im taking a test

OpenStudy (anonymous):

Angles S and V are congruent.  We are given that the triangle are similar, meaning that their sides are proportional. Given this, no matter what proportionality factor, the angles are always the same.

OpenStudy (anonymous):

thanks @CShrix

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Are the following figures similar? Rectangles ABCD and EFGH are shown. AB equals 5. BC equals 25. EF equals 3. FG equals 15.

OpenStudy (anonymous):

can you help me here too?

OpenStudy (anonymous):

@CShrix @happy_to_help

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

@CShrix Whered u go? plz help

OpenStudy (anonymous):

To find congruence, we need to see if the sides are proportional. Let's look at the top side: \[\frac{25}{15}=\frac{5}{3}\] Left side:\[\frac{5}{3}=\frac{5}{3}\] Therefore, given that the respective sides of the rectangles are congruent with the same constant of proportionality, then the rectangles are similar.

OpenStudy (anonymous):

Yes; the corresponding angles are congruent No; the corresponding angles are not congruent Yes; the corresponding sides are proportional No; the corresponding sides are not proportional

OpenStudy (anonymous):

so its yes proportional?

OpenStudy (anonymous):

It is paramount to know the difference between congruence and equality. These rectangles are not equal (=). They have different dimensions and different areas. However, given that they are proportional, then they are considered congruent or similar (≅)

OpenStudy (anonymous):

yes congruent?

OpenStudy (anonymous):

i suck at all geometry

OpenStudy (anonymous):

Technically, both are true. But all rectangles have 4 right angles. But if we look at any two rectangles with 4 right angles, then they might not be similar. In this case, we proved that the rectangles are similar through the proportionality of their sides lengths. I would pick C.

OpenStudy (anonymous):

Thanks :)